# Monopole Acceleration in Intergalactic Magnetic Fields

### v1

@@ 1 Introduction
MMs are accelerated by the cosmic magnetic fields that exist in the universe on various scales – from planets and stars to galaxies, galaxy clusters, filaments, and beyond. Of particular relevance for a cosmic population of MMs are the intergalactic magnetic fields (IGMFs) which permeate the immense intergalactic voids over cosmological distances. The IGMF has been for long constrained by upper limits on the intensity, mainly from cosmological observations of the cosmic microwave background as $B\lesssim 10^{-9}\,\mathrm{G}$. However in the past decade, the existence of IGMF has been suggested from gamma-ray experiments with the non-observation of secondary effects such as gamma-ray cascades, delayed or diffuse emission around blazars [6, 7, 8, 9]. These studies set lower limits on the IGMF strength of $B\gtrsim 10^{-15}\,\mathrm{G}$, when the correlation length $\lambda$ is of Mpc scale or larger. If $\lambda$ is much smaller than a Mpc, the lower limit improves as $\lambda^{-1/2}$. See e.g. [10, 11] for reviews on IGMFs. [^1] Explanations of the blazar observations that do not invoke IGMFs have also been proposed, see e.g. [12, 13].
@@
In light of the observational results, in this paper we analyze in detail the acceleration of MMs in IGMFs. Although IGMFs are extremely weak compared with Galactic fields which are of $B\sim 10^{-6}\,\mathrm{G}$, their very large coherence lengths contribute significantly to the MM acceleration. We demonstrate that IGMFs can easily accelerate the MMs to relativistic velocities. (See also the earlier works [14, 15, 16] which discussed relativistic MMs from Galactic/intergalactic fields.) Moreover, we show that the backreaction of the MMs on the IGMFs can drastically affect the acceleration process and must be taken into account for an accurate calculation of the MM velocity.
@@
Cosmic magnetic fields lose energy when they accelerate MMs. Therefore, the requirement for the survival of the fields imposes an upper limit on the abundance of MMs. This idea was first introduced by Parker, who calculated an upper limit on the MM flux within our Galaxy based on the survival of Galactic magnetic fields [17, 18]. This limit, known as the Parker bound, was later extended to consider the survival of the seed magnetic field of our Galaxy [19], as well as magnetic fields in galactic clusters [20]. Additionally, if IGMFs have a primordial origin, as suggested by numerous studies (see [21, 22] for reviews), their survival also leads to Parker-type constraints [23, 24, 25]. Here we note that in literature, Parker bounds based on Galactic magnetic fields have been derived under the assumption that MMs have a velocity of $v\sim 10^{-3}$, which corresponds to the virial velocity or the peculiar velocity of the Milky Way [26]. However, the bounds strongly depend on the kinetic energy of the MMs, and thus are significantly affected by MM-accelerating processes. As an application of our study of the MM acceleration in IGMFs, we show how Galactic Parker bounds are modified by taking into account the acceleration effects. In particular, we find that the bounds from the survival of seed Galactic magnetic fields can be significantly relaxed in the presence of IGMFs within the observational limits.
@@
This paper is organized as follows. In Section 2 we analyze MM acceleration in IGMFs, and the backreaction on the IGMFs. In Section 3 we show how Galactic Parker bounds are affected by the MM acceleration in IGMFs. We then conclude in Section 4. In Appendix A we evaluate the energy loss of fast-moving MMs in the intergalactic space and inside the Galaxy. In Appendix B we study energy oscillations between MMs and magnetic fields. In Appendix C we compute the MM velocity at Earth by further taking into account the acceleration in Galactic magnetic fields.
@@ 2 Acceleration of monopoles in intergalactic magnetic fields
In this section we derive the velocity of MMs after they have been accelerated in IGMFs for a Hubble time. We start by treating the IGMFs as a background. Then we discuss the backreaction of MMs on the IGMFs, and evaluate the actual velocity that MMs obtain in the intergalactic space.
@@ 2.1 Acceleration in an intergalactic magnetic field background
We model IGMFs with coherence length $\lambda_{\mathrm{I}}$, by dividing the universe into cells of uniform field, with each cell having a size $\lambda_{\mathrm{I}}$. We take the field strength in all cells to have the same value $B_{\mathrm{I}}$, and compute the MM velocity after a Hubble time. For the analyses in this subsection, we ignore numerical factors of order unity.
@@ 2.1.2 Second cell onward
One sees from (2.5) that from the second cell onward, the MM velocity does not change by more than an order-unity factor in each cell. Hence the crossing time for the $N$th cell can be estimated as
@@ 2.1.3 Velocity after a Hubble time
We are now ready to evaluate the velocity of monopoles that have been accelerated in an IGMF background for [^3] As we do not consider time scales larger than one Hubble time, we ignore cosmic expansion. a Hubble time $1/H_{0}$.
@@
In the above discussions we have ignored the energy loss of MMs due to interactions with the intergalactic medium and radiative emissions. In Appendix A we show that these effects are actually negligible.
@@ 2.2 Backreaction on intergalactic magnetic fields
If the velocity $v_{0}$ given in Eqs. (2.10) or (2.11) is much smaller than $v_{\mathrm{max}}$, the backreaction on the IGMFs is negligible. On the other hand if the MM velocity approaches $v_{\mathrm{max}}$, then it implies that the energy of the IGMFs has been transferred to the MMs. However the MMs do not effectively dissipate the gained energy in the intergalactic voids, and they eventually return the energy to the IGMFs. This initiates an energy oscillation between the IGMFs and the population of MMs [23], with the oscillation-averaged MM velocity being of order $v_{\mathrm{max}}$. (We expect the IGMF-MM oscillation to avoid Landau damping; see Appendix B for discussions on this point.) Therefore, the MM velocity in the two cases can collectively be written as
@@
In the case of $v_{0}>v_{\mathrm{max}}$, the expression (2.13) for the maximal velocity can be rewritten in terms of the flux as
@@
In Figure 1 we plot $F_{\mathrm{BR}}$ as a function of the MM mass. The IGMF amplitude is varied as $B_{\mathrm{I}}=10^{-9}\,\mathrm{G}$ (blue), $10^{-12}\,\mathrm{G}$ (purple), $10^{-15}\,\mathrm{G}$ (red), with the correlation length taken as $\lambda_{\mathrm{I}}>1/H_{0}$ (solid lines) and $\lambda_{\mathrm{I}}=1\,\mathrm{Mpc}$ (dashed lines). The solid and dashed lines with the same color join at large masses. The charge is fixed to $g=g_{\mathrm{D}}$. As one goes toward smaller masses each curve drops by a factor $2$; this corresponds to the difference between the nonrelativistic and ultrarelativistic regimes as shown in Eq. (2.16). The displayed results match with the approximate expressions (2.17) and (2.18) up to order-unity factors. One sees that $F_{\mathrm{BR}}$ increases for larger $B_{\mathrm{I}}$, since the IGMF becomes more resilient to the backreaction from the MMs. $F_{\mathrm{BR}}$ at small masses increases also for smaller $\lambda_{\mathrm{I}}$, since the MM acceleration becomes less effective.
(a) $B_{\mathrm{I}}=10^{-15}\,\mathrm{G}$, $\lambda_{\mathrm{I}}=1\,\mathrm{Mpc}$.
(b) $B_{\mathrm{I}}=10^{-15}\,G$, $\lambda_{\mathrm{I}}>1/H_{0}$.
(c) $B_{\mathrm{I}}=10^{-10}\,G$, $\lambda_{\mathrm{I}}=1\,\mathrm{Mpc}$.
(d) $B_{\mathrm{I}}=10^{-10}\,G$, $\lambda_{\mathrm{I}}>1/H_{0}$.
@@ 3.1 Parker bound
Here $\gamma_{\mathrm{MW}}$ is the Lorentz factor of the MMs in the rest frame of the Milky Way, $\lambda_{\mathrm{G}}$ is the coherence length of the Galactic magnetic field, $R$ is the size of the magnetic field region of the Galaxy, $\tau_{\mathrm{gen}}$ is the dynamo time scale for the amplification of the Galactic fields, and $B_{\mathrm{G}}$ is the mean Galactic field strength. [^5] The parameters including $\gamma_{\mathrm{MW}}$ need to satisfy a few conditions for the Parker bound to hold; see Eqs. (2.16)-(2.18) in [25]. These are satisfied for the parameter choices in this paper. One can further require the survival of the initial seed magnetic field of the Galaxy, the flux bound from which takes the same form except for that $B_{\mathrm{G}}$ is taken as the seed field strength [19]. [^6] We note that the seed Parker bound of the form (3.1) assumes a seed field in the recent universe. The bound can be modified for seed fields at high redshifts.
@@ 3.3 Parker bound in the presence of intergalactic magnetic fields
In Figure 4 we plot the Parker bound in Eq. (3.1), with the MM Lorentz factor $\gamma_{\mathrm{MW}}$ evaluated through Eq. (3.2). The MM charge is fixed to $g=g_{\mathrm{D}}$. The gray solid line shows the bound from the survival of Galactic fields with $B_{\mathrm{G}}=10^{-6}\,\mathrm{G}$, while the dashed curves are from seed fields with an assumed strength of $B_{\mathrm{G}}=10^{-11}\,\mathrm{G}$. The other Galactic parameters are fixed to $\lambda_{\mathrm{G}}=1\,\mathrm{kpc}$, $R=10\,\mathrm{kpc}$, $\tau_{\mathrm{gen}}=10^{8}\,\mathrm{yr}$, and $v_{\mathrm{p}}=10^{-3}$. The IGMF strength is varied between $B_{\mathrm{I}}=10^{-9}\,\mathrm{G}$ and $B_{\mathrm{I}}=10^{-13}\,\mathrm{G}$. IGMFs within this range do not affect the Galactic bound, which hence is shown by a single solid line. On the other hand, the seed bound is strongly affected by the presence of IGMFs; dashed lines of different colors correspond to different IGMF strength as shown in the plot legend. The Parker bounds for a fixed MM velocity of $v_{\mathrm{MW}}=10^{-3}$, as have been assumed in the literature, are reproduced for sufficiently weak IGMFs. The Galactic and seed bounds in the weak IGMF limit are shown respectively by the gray solid and red dashed curves. All the curves join at large masses, in other words, IGMFs modify the seed bound at intermediate to low masses. We also note that with the choice of parameters, the flux bounds are independent of the IGMF correlation length, as long as it is as large as $\lambda_{\mathrm{I}}\gtrsim 10^{-5}\,\mathrm{Mpc}$. This is because for values of $F_{\mathrm{MW}}$ that saturate the bound, $v_{\mathrm{MW}}$ is set either by $v_{\mathrm{p}}$ or $v_{\mathrm{max}}$, both of which do not depend on $\lambda_{\mathrm{I}}$. This can also be seen from Figures 4 and 1 showing that the upper bounds of $F_{\mathrm{MW}}$ are larger than the threshold value $F_{\mathrm{BR}}$ for the backreaction to the IGMFs to become relevant.
@@
For the choice of parameters in the plot, the upper bound on the flux from Galactic fields ($B_{G}=10^{-6}\,\mathrm{G}$) in the weak IGMF limit is of $F_{\mathrm{MW}}\sim 10^{-16}\,\mathrm{cm}^{-2}\mathrm{sec}^{-1}\mathrm{sr}^{-1}$ or larger. The threshold kinetic energy between the two expressions for the Parker bound is $m(\gamma_{\mathrm{MW}}-1)\sim 10^{11}\,\mathrm{GeV}$. For flux values saturating the bound, the energy that MMs acquire from IGMFs with $B_{I}\leq 10^{-9}\,\mathrm{G}$ is smaller than the threshold, which is why the Galactic bound is not altered by IGMFs. On the other hand, seed fields ($B_{G}=10^{-11}\,\mathrm{G}$) give an upper bound in the weak IGMF limit of $F_{\mathrm{MW}}\sim 10^{-21}\,\mathrm{cm}^{-2}\mathrm{sec}^{-1}\mathrm{sr}^{-1}$ or larger, with the threshold energy between the two expressions being $m(\gamma_{\mathrm{MW}}-1)\sim 10^{6}\,\mathrm{GeV}$. For the flux values, the energy can exceed the threshold with IGMFs of $B_{\mathrm{I}}\gtrsim 10^{-12}\,\mathrm{G}$, hence the seed Parker bound is modified.
In order to understand the behavior of the modified Parker bound, note that substituting Eq. (3.2) for $\gamma_{\mathrm{MW}}$ into the first line of Eq. (3.1) yields an upper limit which itself depends on the MM flux. In particular when $v_{\mathrm{MW}}=v_{\mathrm{max}}$ (i.e. $v_{\mathrm{p}}<v_{\mathrm{max}}<v_{0}$), one sees from Eq. (2.16) that $(\gamma v)_{\mathrm{MW}}\propto 1/\left(mF_{\mathrm{MW}}\right)$. For sufficiently small masses such that $v_{\mathrm{MW}}$ is relativistic, the kinetic energy scales as $m\gamma_{\mathrm{MW}}\propto 1/F_{\mathrm{MW}}$. Thus the first line of Eq. (3.1) yields an inequality of the form $F_{\mathrm{MW}}<\kappa/F_{\mathrm{MW}}$ where $\kappa$ is a factor that is independent of the MM mass and abundance. Solving this gives a flux bound $F_{\mathrm{MW}}<\kappa^{1/2}$, which is independent of the MM mass, as shown in the plot for the modified seed Parker bound at small masses. On the other hand for larger masses with nonrelativistic $v_{\mathrm{MW}}$, the kinetic energy scales as $mv_{\mathrm{MW}}^{2}/2\propto 1/mF_{\mathrm{MW}}^{2}$, and hence the first line of Eq. (3.1) becomes $F_{\mathrm{MW}}<\kappa/mF_{\mathrm{MW}}^{2}$. This gives a bound $F_{\mathrm{MW}}<(\kappa/m)^{1/3}$, which is stronger for larger masses, as is seen for the middle parts of the modified seed bounds in the plot.

### v2

@@ 1 Introduction
MMs are accelerated by the cosmic magnetic fields that exist in the universe on various scales – from planets and stars to galaxies, galaxy clusters, filaments, and beyond. Of particular relevance for a cosmic population of MMs are the intergalactic magnetic fields (IGMFs) which permeate the immense intergalactic voids over cosmological distances. The IGMF has been for long constrained by upper limits on the intensity, mainly from cosmological observations of the cosmic microwave background as $B\lesssim 10^{-9}\,\mathrm{G}$. However in the past decade, the existence of IGMF has been suggested from gamma-ray experiments with the non-observation of extended halos or delayed emission around blazars formed by secondary photons [6, 7, 8, 9]. These studies set lower limits on the IGMF strength of $B\gtrsim 10^{-15}\,\mathrm{G}$, if the correlation length $\lambda$ is of Mpc scale or larger. If $\lambda$ is much smaller than a Mpc, the lower limit further improves as $\lambda^{-1/2}$. See e.g. [10, 11] for reviews on IGMFs. [^1] A non-IGMF explanation of the blazar observations that invokes a beam plasma instability was also proposed [12, 13], although this mechanism has been questioned in, e.g., [14, 15].
@@
In light of the observational results, in this paper we analyze in detail the acceleration of MMs in IGMFs. Although IGMFs are extremely weak compared with Galactic fields which are of $B\sim 10^{-6}\,\mathrm{G}$, their very large coherence lengths contribute significantly to the MM acceleration. We demonstrate that IGMFs can easily accelerate the MMs to relativistic velocities. (See also the earlier works [16, 17, 18] which discussed relativistic MMs from Galactic/intergalactic fields.) Moreover, we show that the backreaction of the MMs on the IGMFs can drastically affect the acceleration process and must be taken into account for an accurate calculation of the MM velocity. As we are interested in the average velocity of the cosmic population of MMs, we mainly focus on the universe-filling IGMFs. There can be additional extragalactic magnetic fields, for instance fields in cosmic filaments, as well as fields transported by galactic winds (see, e.g. [19, 20, 21, 22, 23, 24]). However it is expected that such fields do not fill the entire universe [25], and thus can only affect the velocity of a small fraction of the cosmic MMs. We hence exclude them from our analysis.
@@
Cosmic magnetic fields lose energy when they accelerate MMs. Therefore, the requirement for the survival of the fields imposes an upper limit on the abundance of MMs. This idea was first introduced by Parker, who calculated an upper limit on the MM flux within our Galaxy based on the survival of Galactic magnetic fields [26, 27]. This limit, known as the Parker bound, was later extended to the survival of the seed magnetic field of our Galaxy [28], as well as magnetic fields in galactic clusters [29]. Additionally, if IGMFs have a primordial origin, as suggested by numerous studies (see [30, 31] for reviews), their survival also leads to Parker-type constraints [32, 33, 34]. Here we note that in literature, Parker bounds based on Galactic magnetic fields have been derived under the assumption that MMs have a velocity of $v\sim 10^{-3}$, which corresponds to the virial velocity or the peculiar velocity of the Milky Way Galaxy [35]. However, the bounds strongly depend on the kinetic energy of the MMs, and thus are significantly affected by processes that accelerate MMs before they enter the Galaxy. As an application of our study of the MM acceleration in IGMFs, we show how Galactic Parker bounds are modified by taking into account the acceleration effects. In particular, we find that the bounds from the survival of seed Galactic magnetic fields can be significantly relaxed in the presence of IGMFs within the observational limits.
@@
This paper is organized as follows. In Section 2 we analyze MM acceleration in IGMFs, and the backreaction on the IGMFs. In Section 3 we show how Galactic Parker bounds are affected by the MM acceleration in IGMFs. We then conclude in Section 4. In Appendix A, we consider the hypothesis that the IGMFs have a primordial origin and show that MM acceleration in primordial fields does not modify the order-of-magnitude results of the main text. In Appendix B we evaluate the energy loss of fast-moving MMs in the intergalactic space and inside the Galaxy. In Appendix C we study energy oscillations between MMs and magnetic fields. In Appendix D we compute the MM velocity at Earth by further taking into account the acceleration in Galactic magnetic fields.
@@ 2 Acceleration of monopoles in intergalactic magnetic fields
In this section, we derive the velocity of MMs after they have been accelerated in IGMFs for a Hubble time. We start by treating the IGMFs as a background. Then we discuss the backreaction of MMs on the IGMFs, and evaluate the actual velocity that MMs obtain in the intergalactic space.
@@ 2.1 Acceleration in an intergalactic magnetic field background
Let us model IGMFs with coherence length $\lambda_{\mathrm{I}}$, by dividing the universe into cells of uniform field, with each cell having a size $\lambda_{\mathrm{I}}$. We take the field strength in all cells to have the same value $B_{\mathrm{I}}$, and compute the MM velocity after a Hubble time. We will occasionally ignore numerical factors of order unity. [^2] The final results of this paper are only mildly sensitive to the ignored factors, such that the results give the correct order of magnitude (unless there is a drastic modification to the magnetic field model itself.) The plots can also be trusted at the order-of-magnitude level.
@@ 2.1.2 Second cell onward
One sees from (2.5) that from the second cell onward, the MM velocity does not change in each cell by more than an order-unity factor. Hence the crossing time for the $N$th cell can be estimated as
@@ 2.1.3 Velocity after a Hubble time
We are now ready to evaluate the velocity of monopoles that have been accelerated in an IGMF background. We consider the acceleration over a Hubble time $1/H_{0}$, and hence ignore cosmic expansion. In Appendix A we show that, even if the IGMF are remnants of primordial fields and the MMs have been accelerated since the early universe, our estimate of the final MM velocity is modified only by order-unity factors.
@@
In the above discussions we have ignored the energy loss of MMs due to interactions with the intergalactic medium and radiative emissions. In Appendix B we show that these effects are actually negligible.
@@ 2.2 Backreaction on intergalactic magnetic fields
If the velocity $v_{0}$ given in Eqs. (2.10) or (2.11) is much smaller than $v_{\mathrm{max}}$, the backreaction on the IGMFs is negligible. On the other hand if the MM velocity approaches $v_{\mathrm{max}}$, then it implies that the energy of the IGMFs has been transferred to the MMs. However the MMs do not effectively dissipate the gained energy into the intergalactic medium as shown in Appendix B, and thus they eventually return the energy to the IGMFs. This initiates an energy oscillation between the IGMFs and the population of MMs [32], with the oscillation-averaged MM velocity being of order $v_{\mathrm{max}}$. We expect the IGMF-MM oscillation to avoid Landau damping; see Appendix C for discussions on this point.
The MM velocity in the above two cases can collectively be written as
@@
In the case where $v_{0}>v_{\mathrm{max}}$, the expression (2.13) for the maximal velocity can be rewritten in terms of the flux as
@@
(a) $B_{\mathrm{I}}=10^{-15}\,\mathrm{G}$, $\lambda_{\mathrm{I}}=1\,\mathrm{Mpc}$.
(b) $B_{\mathrm{I}}=10^{-15}\,G$, $\lambda_{\mathrm{I}}>1/H_{0}$.
(c) $B_{\mathrm{I}}=10^{-10}\,G$, $\lambda_{\mathrm{I}}=1\,\mathrm{Mpc}$.
(d) $B_{\mathrm{I}}=10^{-10}\,G$, $\lambda_{\mathrm{I}}>1/H_{0}$.
@@
In Figure 1 we plot $F_{\mathrm{BR}}$ as a function of the MM mass. The IGMF amplitude is varied as $B_{\mathrm{I}}=10^{-9}\,\mathrm{G}$ (blue), $10^{-12}\,\mathrm{G}$ (purple), $10^{-15}\,\mathrm{G}$ (red), with the correlation length taken as $\lambda_{\mathrm{I}}>1/H_{0}$ (solid lines) and $\lambda_{\mathrm{I}}=1\,\mathrm{Mpc}$ (dashed lines). The solid and dashed lines with the same color join at large masses. The charge is fixed to $g=g_{\mathrm{D}}$. We show MM masses down to $1\,\mathrm{GeV}$, however we remark that this value is a few orders of magnitude smaller than the latest experimental lower limit [36]. As one goes toward smaller masses each curve drops by a factor $2$; this corresponds to the difference between the nonrelativistic and ultrarelativistic regimes as shown in Eq. (2.16). The displayed results match with the approximate expressions (2.17) and (2.18) up to order-unity factors. One sees that $F_{\mathrm{BR}}$ increases for larger $B_{\mathrm{I}}$, since the IGMF becomes more resilient to the backreaction from the MMs. $F_{\mathrm{BR}}$ at small masses increases also for smaller $\lambda_{\mathrm{I}}$, since the MM acceleration becomes less effective.
@@ 3.1 Parker bound
Here $\gamma_{\mathrm{MW}}$ is the Lorentz factor of the MMs in the rest frame of the Milky Way, $\lambda_{\mathrm{G}}$ is the coherence length of the Galactic magnetic field, $R$ is the size of the magnetic field region of the Galaxy, $\tau_{\mathrm{gen}}$ is the dynamo time scale for the amplification of the Galactic fields, and $B_{\mathrm{G}}$ is the mean Galactic field strength. [^5] The parameters including $\gamma_{\mathrm{MW}}$ need to satisfy a few conditions for the Parker bound to hold; see Eqs. (2.16)-(2.18) in [34]. These are satisfied for the parameter choices in this paper. One can further require the survival of the initial seed magnetic field of the Galaxy, the flux bound from which takes the same form except for that $B_{\mathrm{G}}$ is taken as the seed field strength [28]. [^6] We note that the seed Parker bound of the form (3.1) neglects the cosmic expansion since the initial time when the dynamo begins to operate. The bound can be modified for seed fields at high redshifts.
@@ 3.3 Parker bound in the presence of intergalactic magnetic fields
In Figure 4 we plot the Parker bound in Eq. (3.1), with the MM Lorentz factor $\gamma_{\mathrm{MW}}$ evaluated through Eq. (3.2). The MM charge is fixed to $g=g_{\mathrm{D}}$. The gray solid line shows the bound from the survival of Galactic fields with $B_{\mathrm{G}}=10^{-6}\,\mathrm{G}$, while the dashed curves are from seed fields with an assumed strength of $B_{\mathrm{G}}=10^{-11}\,\mathrm{G}$. The other Galactic parameters are fixed to $\lambda_{\mathrm{G}}=1\,\mathrm{kpc}$, $R=10\,\mathrm{kpc}$, $\tau_{\mathrm{gen}}=10^{8}\,\mathrm{yr}$, and $v_{\mathrm{p}}=10^{-3}$. The IGMF strength is varied between $B_{\mathrm{I}}=10^{-9}\,\mathrm{G}$ and $B_{\mathrm{I}}=10^{-13}\,\mathrm{G}$. IGMFs within this range do not affect the Galactic bound, which hence is shown by a single solid line. On the other hand, the seed bound is strongly affected by the presence of IGMFs; dashed lines of different colors correspond to different IGMF strengths as shown in the plot legend. The Parker bounds for a fixed MM velocity of $v_{\mathrm{MW}}=10^{-3}$, as have been assumed in the literature, are reproduced for sufficiently weak IGMFs. The Galactic and seed bounds in the weak IGMF limit are shown respectively by the gray solid and red dashed curves. All the curves join at large masses, in other words, IGMFs modify the seed bound at intermediate to low masses. We also note that with the choice of parameters, the flux bounds are independent of the IGMF correlation length, as long as it is as large as $\lambda_{\mathrm{I}}\gtrsim 10^{-5}\,\mathrm{Mpc}$. This is because for values of $F_{\mathrm{MW}}$ that saturate the bound, $v_{\mathrm{MW}}$ is set either by $v_{\mathrm{p}}$ or $v_{\mathrm{max}}$, both of which do not depend on $\lambda_{\mathrm{I}}$. This can also be seen from Figures 4 and 1 showing that the upper bounds of $F_{\mathrm{MW}}$ are larger than the threshold value $F_{\mathrm{BR}}$ for the backreaction to the IGMFs to become relevant.
@@
For the choice of parameters in the plot, the upper bound on the flux from Galactic fields ($B_{G}=10^{-6}\,\mathrm{G}$) in the weak IGMF limit is of $F_{\mathrm{MW}}\sim 10^{-16}\,\mathrm{cm}^{-2}\mathrm{sec}^{-1}\mathrm{sr}^{-1}$ or larger. The threshold kinetic energy between the two expressions for the Parker bound is $m(\gamma_{\mathrm{MW}}-1)\sim 10^{11}\,\mathrm{GeV}$. For flux values saturating the bound, the energy that MMs acquire from IGMFs with $B_{I}\leq 10^{-9}\,\mathrm{G}$ is smaller than the threshold, which is why the Galactic bound is not altered by IGMFs. On the other hand, seed fields ($B_{G}=10^{-11}\,\mathrm{G}$) give an upper bound in the weak IGMF limit of $F_{\mathrm{MW}}\sim 10^{-21}\,\mathrm{cm}^{-2}\mathrm{sec}^{-1}\mathrm{sr}^{-1}$ or larger, with the threshold energy between the two expressions being $m(\gamma_{\mathrm{MW}}-1)\sim 10^{6}\,\mathrm{GeV}$. For these flux values, the energy can exceed the threshold with IGMFs of $B_{\mathrm{I}}\gtrsim 10^{-12}\,\mathrm{G}$, hence the seed Parker bound is modified.
In order to understand the behavior of the modified Parker bound, note that substituting Eq. (3.2) for $\gamma_{\mathrm{MW}}$ into the first line of Eq. (3.1) yields an upper limit which itself depends on the MM flux. In particular when $v_{\mathrm{MW}}=v_{\mathrm{max}}$ (i.e. $v_{\mathrm{p}}<v_{\mathrm{max}}<v_{0}$), one sees from Eq. (2.16) that $(\gamma v)_{\mathrm{MW}}\propto 1/\left(mF_{\mathrm{MW}}\right)$. For sufficiently small masses such that $v_{\mathrm{MW}}$ is relativistic, the kinetic energy scales as $m\gamma_{\mathrm{MW}}\propto 1/F_{\mathrm{MW}}$. Thus the first line of Eq. (3.1) yields an inequality of the form $F_{\mathrm{MW}}<\kappa/F_{\mathrm{MW}}$ where $\kappa$ is a factor that is independent of the MM mass and abundance. Solving this gives a flux bound $F_{\mathrm{MW}}<\kappa^{1/2}$, which is independent of the MM mass, as shown in the plot for the modified seed Parker bound at small masses. On the other hand for larger masses with nonrelativistic $v_{\mathrm{MW}}$, the kinetic energy scales as $mv_{\mathrm{MW}}^{2}/2\propto 1/mF_{\mathrm{MW}}^{2}$, and hence the first line of Eq. (3.1) takes a form $F_{\mathrm{MW}}<\kappa/mF_{\mathrm{MW}}^{2}$. This gives a bound $F_{\mathrm{MW}}<(\kappa/m)^{1/3}$, which is stronger for larger masses, as is seen for the middle parts of the modified seed bounds in the plot.